when should I use 10 ^(2pH) in equilibrium concentration questions?
1
Expert's answer
2020-05-11T14:24:25-0400

According to the definition, pH is the decimal logarithm of the reciprocal hydrogen ion concentration in a solution.

As a result, 10-pH represents the concentration of the hydrogen ions.

According to the common reaction:

AH = A- + H+

From here the equilibrium constant is:

Keq = [A-] [H+] / [AH]

If AH is a strong acid, it completely dissociates into [A-] and [H+]. As a result, [A-]=[H+]. From here:

Keq = [H+]2 / [AH]

As [H+] = 10-pH, [H+]2 = 10-2pH.

As a result, 10-2pH should be used if we have a strong acid.